need intercepting points for these two equations: f(x)=-2/315x^2+630 and 4/3x+420
thnx!

Answer by stanbon(75887)

need intercepting points for these two equations:
f(x)=-2/315x^2+630 and 4/3x+420
--------
Solve::
(-2/315)x^2 + 630 = (4/3)x + 420
-----
(-2/315)x^2 - (4/3)x + 210 = 0
-------
I graphed it and found these Real Number solutions::
x = 105 and x = -315
========================
When x = 105, y = (-2/315)105^2 + 630 = 560
When x = -315, y = (-2/315)315^2 + 630 = 0
-------------
Cheers,
Stan H.
-----------------------